Hi. Can I have some help with functions?
Evaluate the function at the indicated values.
h(t)=t+1/t;
h(1),h(-1),h(2),h(1/2),h(x),h(1/x)

Answer by ilana(307)

You certainly can. When you change h(t) to h(1), you are replacing every t with 1. So h(1)=(1)+1/(1). Since division comes before addition in the order of operations, do the division first (assuming you typed this in exactly how it appears). So it is 1+1=2. So h(1)=2.
h(-1)=(-1)+1/(-1)=-1+-1=-2
Now I'll skip the middle ones for you to do, but I'll do the last since it is most tricky. Try it yourself and check your answer after.
h(1/x)=(1/x)+1/(1/x) = (1/x)+x